OpenStudy (anonymous):

Sqrt(5x^2)×sqrt(40x)

OpenStudy (tkhunny):

Find the prime factorizations: \(\sqrt{5*40*x^2*x}\)

OpenStudy (anonymous):

\[=\sqrt(5)*\sqrt(x^2)*\sqrt(40)*\sqrt(x) \] Combine constants multiplicatively \[ =\sqrt(200)*x*\sqrt(x) \] combine powers of x (\sqrt = ^1/2) and extract factor of 4 from 200 [= 4*50] \[ =2*\sqrt(50)*x^{3/2} \] 50 = 2*25 = 2*(5^2), so extract 5 \[= 10*\sqrt(2)*x^{3/2}\] \[x\]

OpenStudy (tkhunny):

Look for factors,don't just multiply willy-nilly. \(\sqrt{5*40*x^{2}*x}\) We know that x >= 0. \(x\sqrt{5*40*x}\) There's a factor of 5 in 40. \(5x\sqrt{8*x}\) There's a factor of 4 in 8. \(10x\sqrt{2*x}\) And it's done.